OpenStudy (anonymous):

Two chords intersect inside a circle. The lengths of the segments of one chord are 4 and 6. The lengths of the segments of the other chord are 3 and 7, 8, OR 9?!?!?!!?!?!?!?!?!!?!?!?!?!?!?!!?!?!?!?!?!?!?!?!?

Directrix (directrix):

Read the theorem about two chords intersecting in a circle. Then, we have to set up the equation.

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

3/24

Directrix (directrix):

4*6 = 3*x 3x = 24 x = 8 ---> Is that what we said is correct? @autumnpoteet